# A Systems Thinking Approach to Algorithmic Fairness

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; 2025; Athens, Greeceisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Causal reasoning and diagnosticsccs: Computing methodologies Probabilistic reasoningccs: Computing methodologies Reasoning about belief and knowledgeccs: Applied computing Lawccs: Applied computing Sociologyccs: Applied computing Economicsccs: Applied computing Psychologyccs: Social and professional topics Governmental regulations
@@ 1. Introduction
To address the algorithmic fairness problem, we need to take a sociotechnical approach. We will begin by analyzing the social aspects of this problem using first principles. We will explore the nature of free will to develop an understanding of moral responsibility. We will review the structure versus agency debate in sociology and attribution theory in psychology to address the controversial question: Why do we believe disparities exist across protected groups like race or gender? This will give us insights into the sources of bias in the data generating process, showing us what different interventions are needed to make a decision fair. We can then use this knowledge to build a better understanding of the political debate around fairness, enabling us to develop a deeper and more holistic way to analyze this complex and difficult problem.
@@
Finally, we will bring together the social and technical aspects of the fairness problem. We will build a system map to show how the social sciences and computer science are related to each other. We will show the dualistic nature of the fairness problem across the social sciences and how this relates to Pearl’s two fundamental laws of causal inference. Hopefully, this will help us to finally bridge the sociotechnical fairness gap.
@@ 2. Related work
Another key difference is that their approaches relid solely on causal inference, which has limitations in its ability to model the data generating process. Our approach leverages systems thinking and specifically system dynamics to gain a fuller understanding of the data generating process through the modeling feedback loops. Furthermore, we link supervised machine learning to causal inference by recognizing that a CBN can be used as a higher level abstraction of a black box ML model. By bringing machine learning, causal inference, and systems thinking together, we can model the different aspects of the algorithmic fairness problem.
@@ 6.1. Case study: Racial disparities in the US
We can use the CBNs from Figure 2 and the CLD from Figure 6 to model the complex political debate over racial disparities in the US. According to the US Census Bureau, median household income in 2021 was \$100,573 for Asians, \$74,932 for Whites, \$57.671 for Hispanics, and \$48,297 for Blacks(Survey, 2021). Meanwhile, median household wealth in 2021 was \$320,900 for Asians, \$250,400 for Whites, \$48,700 for Hispanics, and \$27,100 for Blacks(Kochnar and Moslimani, 2023). We would like to understand why Whites and Asians have higher median household income and wealth than Blacks and Hispanics.
@@
Placing Blacks and Hispanics into higher level positions than is warranted by their merits and qualifications could also exact a harmful psychological toll. This could reinforce their views that they are victims of an unjust system, leading to a sense of learned helplessness and a loss of agency. This could result in reduced productivity, which is the primary determinant of income and wealth in the capitalist economic system of the US. Rather than helping Blacks and Hispanics become more competitive against Whites and Asians, affirmative action may have instead made them less competitive.
@@
Affirmative action is a low-leverage solution that should only be used to modestly decrease racial disparities. It is not a fundamental, high-leverage solution for reducing racial disparities over the long-term. As long as there are racial disparities in the mediator $W$, then there will be racial disparities in the outcome $Y$. This would mean that racial disparities in household income and wealth in the US would continue to persist. The focus should instead be on closing disparities in the mediator $W$. For example, Democrats could call for increased government funding of education and training programs. On the other hand, Republicans could call for the banning of racial preferences in high-stakes decisions.
Hopefully, these models provide a basis for arguing that while historical policies like slavery and Jim Crow generated the initial racial disparities between Whites and Blacks, contemporary policies like affirmative action and DEI may have reinforced those disparities. Affirmative action should only be used as a temporary solution for marginally closing the racial equity gap and if it is used, then there should be a long-term plan to eventually switch back to fairness through unawareness. If an addictive dependency on affirmative action develops that diverts attention away from internal solutions for closing the merit and qualifications gap, then the affirmative action policy should be ended. Systems thinking shows us that AI researchers should not be so quick to dismiss fairness through unawareness as a fundamental solution for closing the racial equity gap.
@@ 8. Modeling fairness as a systems map
As mentioned before, algorithmic fairness is a sociotechnical problem. We need to find a way to bridge the social and technical aspects of the fairness problem. In each of the social science fields, there exists a dichotomy that represents different viewpoints for understanding the fairness problem. This allows us to understand the fundamentally ideological and political debate over fairness, which centers around how we attribute the causes for an outcome. The systems map shows a causal bridge that links this dualism in the social sciences to Pearl’s two fundamental laws of causal inference, which was discussed in section 5.1. Hopefully this systems map provides a very compact and high level representation for how the different fields of fairness relate to each another.
@@ 9. Conclusion
This approach allow us to develop a more unified and holistic view of the fairness and machine learning problem, providing the essential connections between computer science and the law. This is the key to building AI systems that are aligned with society’s democratic values and the policymakers’ political agendas. It will enable businesses and regulators to communicate with each other using a common causal language, one that can be used by both data scientists and lawyers. These models may even aid in AI deregulation by identifying mathematical flaws in the existing laws and regulations that are caused by paradoxes. This will help businesses to comply with antidiscrimination laws and regulations, reducing business uncertainty and expensive litigation while providing a much needed layer of transparency and trust for consumers of AI products and services.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: 10.48550/arXiv.2412.16641conference: ACM Conference on Fairness, Accountability, and Transparency; 2025; Athens, Greeceisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Causal reasoning and diagnosticsccs: Computing methodologies Probabilistic reasoningccs: Computing methodologies Reasoning about belief and knowledgeccs: Applied computing Lawccs: Applied computing Sociologyccs: Applied computing Economicsccs: Applied computing Psychologyccs: Social and professional topics Governmental regulations
@@ 1. Introduction
To address the algorithmic fairness problem, we need to take a sociotechnical approach. We will begin by analyzing the social aspects of this problem using first principles. We will explore the nature of free will to develop an understanding of moral responsibility. We will review the structure versus agency debate in sociology and attribution theory in psychology to address the controversial question: Why do we believe disparities exist across protected groups like race or gender? This will give us insights into the sources of bias in the data generating process, showing us which interventions are needed to make a decision fair. We can then use this knowledge to build a better understanding of the political debate around fairness, enabling us to develop a deeper and more holistic way to analyze this complex and difficult problem.
@@
Finally, we will bring together the social and technical aspects of the fairness problem. We will build a system map to show how the social sciences and computer science are related to each other. We will show the dualistic nature of the fairness problem across the social sciences and how this relates to Pearl’s two fundamental laws of causal inference. This may allow us to finally bridge the sociotechnical fairness gap.
@@ 2. Related work
Another key difference is that their approaches relied solely on causal inference, which has limitations in its ability to model the data generating process. Our approach leverages systems thinking and specifically system dynamics to gain a fuller understanding of the data generating process through the modeling of feedback loops. Furthermore, we link supervised machine learning to causal inference by recognizing that a CBN can be used as a higher level abstraction of a black box ML model. By bringing machine learning, causal inference, and systems thinking together, we can model the different aspects of the algorithmic fairness problem.
@@ 6.1. Case study: Racial disparities in the US
We can use the CBNs from Figure 2 and the CLD from Figure 6 to model the complex political debate over racial disparities in the US. According to the US Census Bureau, median household income in 2021 was \$100,573 for Asians, \$74,932 for Whites, \$57,671 for Hispanics, and \$48,297 for Blacks(Survey, 2021). Meanwhile, median household wealth in 2021 was \$320,900 for Asians, \$250,400 for Whites, \$48,700 for Hispanics, and \$27,100 for Blacks(Kochnar and Moslimani, 2023). We would like to understand why Whites and Asians have higher median household income and wealth than Blacks and Hispanics.
@@
Placing Blacks and Hispanics into higher level positions than is warranted by their merits and qualifications could also exact a harmful psychological toll. This could reinforce their views that they are victims of an unjust system, leading to a sense of learned helplessness and a loss of agency. This could result in a reduction of human capital and productivity, which is the primary determinant of income and wealth in the primarily capitalist economic system of the US(Becker, 1964). Rather than helping Blacks and Hispanics become more competitive against Whites and Asians, affirmative action may have instead made them less competitive.
@@
Affirmative action is a low-leverage solution that should only be used to modestly decrease racial disparities. It is not a fundamental, high-leverage solution for reducing racial disparities over the long-term. As long as there are racial disparities in the mediator $W$, then there will be racial disparities in the outcome $Y$. This would mean that racial disparities in household income and wealth in the US would continue to persist. The focus should instead be on closing disparities in the mediator $W$. For example, Democrats could call for increased government funding of education, training, and health programs to help minorities improve their human capital. On the other hand, Republicans could call for the banning of racial preferences in high-stakes decisions to keep everyone accountable.
These models provide a basis for arguing that while historical policies like slavery and Jim Crow generated the initial racial disparities between Whites and Blacks, contemporary policies like affirmative action and DEI may have reinforced those disparities. Affirmative action should only be used as a temporary solution for marginally closing the racial equity gap and if it is used, then there should be a long-term plan to eventually switch back to fairness through unawareness. If an addictive dependency on affirmative action develops that diverts attention away from internal solutions for closing the merit and qualifications gap, then the affirmative action policy should be ended. Systems thinking shows us that AI researchers should not have been so quick to dismiss fairness through unawareness as a fundamental solution for closing the racial equity gap.
@@ 8. Modeling fairness as a systems map
As mentioned before, algorithmic fairness is a sociotechnical problem. We need to find a way to bridge the social and technical aspects of the fairness problem. In each of the social science fields, there exists a dichotomy that represents different viewpoints for understanding the fairness problem. This allows us to understand the fundamentally ideological and political debate over fairness, which centers around how we attribute the causes for an outcome. The systems map shows a causal bridge that links this dualism in the social sciences to Pearl’s two fundamental laws of causal inference, which was discussed in section 5.1. This systems map can provide a very compact and high level representation for how the different fields of fairness relate to each another.
@@ 9. Conclusion
This approach allow us to develop a more unified and holistic view of the fairness and machine learning problem, providing the essential connections between computer science, politics, and the law. This is the key to building AI systems that are aligned with society’s democratic values and the policymakers’ political agendas. It will enable businesses and regulators to communicate with each other using a common causal language, one that can be used by both data scientists and lawyers. These models may even aid in AI deregulation by identifying mathematical flaws in the existing laws and regulations that are caused by paradoxes. This will help businesses to comply with antidiscrimination laws and regulations, reducing business uncertainty and expensive litigation while providing a much needed layer of transparency and trust for consumers of AI products and services.